### v1

@@ 6.3.1 Feature Importance Rankings
Both methods consistently identified primary outputs (PO), flip-flop outputs (ffo), and logic gate fan-in (LGFi) as the most discriminative features for trojan detection, aligning with domain knowledge that trojans often manifest as structural anomalies in these circuit characteristics.
@@ 7 Conclusion
The comparative evaluation on 11,392 gates from Trust-Hub benchmark circuits yields several results. For detection performance, XGBoost classification with optimized decision threshold achieves 46.15% precision (95% CI: [42.06%, 50.24%]) and 52.17% recall (95% CI: [45.65%, 58.70%]) on severely imbalanced data (275:1 benign-to-trojan ratio), representing a 9-fold improvement in precision over prior work (Hasegawa et al.: 5.13% precision, 82.6% recall). This improvement reduces false positive rates from approximately 5.6% to 0.25%, addressing a limitation for practical deployment where security engineers must manually investigate flagged gates. Statistical validation via McNemar’s test confirms these improvements are statistically significant ($\chi^{2}=1431.25$, $p<0.001$), with the 15.11 percentage point accuracy improvement (99.56% vs. 84.45%) indicating meaningful practical differences beyond statistical significance alone.

### v2

@@ 6.3.1 Feature Importance Rankings
Both methods consistently identified PO, FFo, and LGFi as the most discriminative features for trojan detection, aligning with domain knowledge that trojans often manifest as structural anomalies in these circuit characteristics.
@@ 7 Conclusion
The comparative evaluation on 11,392 gates from the Trust-Hub benchmark dataset yields several results. For detection performance, XGBoost classification with optimized decision threshold achieves 46.15% precision (95% CI: [42.06%, 50.24%]) and 52.17% recall (95% CI: [45.65%, 58.70%]) on severely imbalanced data (275:1 benign-to-trojan ratio), representing a 9-fold improvement in precision over prior work (Hasegawa et al.: 5.13% precision, 82.6% recall). This improvement reduces false positive rates from approximately 5.6% to 0.25%, addressing a limitation for practical deployment where security engineers must manually investigate flagged gates. Statistical validation via McNemar’s test confirms these improvements are statistically significant ($\chi^{2}=1431.25$, $p<0.001$), with the 15.11 percentage point accuracy improvement (99.56% vs. 84.45%) indicating meaningful practical differences beyond statistical significance alone.